The difference of two numbers is 1. 1/7 the larger number less 1/9 the smaller number is 1. Find the numbers.
Answer by oscargut(2103)

Let x the larger number and y the other
then
x-y=1
x/7-y/9=1 (x63)
then
x-y=1 (-7)
9x-7y=63
-7x+7y=-7
9x-7y=63
then (doing the addition)
2x=56
x=28 and y=27